tinagoff:

Match the equation with it's graph -4×-2y=8

563blackghost:

Easier to turn this into a slope-intercept form function so you can graph it. \(\large\bf{-4x-2y=8}\) \(\large\bf{{\color{red}{4x~+}~(-4x)-2y=8~{\color{Red}{+~4x}}}}\) \(\Large\bf{\frac{-2y}{\color{Red}{2}} = \frac{8}{\color{red}{2}} + \frac{4x}{\color{Red}{2}}}\) \(\Large\bf{y=-2x-4}\) Now you can graph it. You can use desmos.com to try and graph :)

563blackghost:

Welcome to QC :)

Falconmaster:

Do you still need help with this?

silvernight269:

https://www.desmos.com/calculator

Vocaloid:

Closed due to answered.